Solve (16^x)(x^x-2)=(1/64)
Answer by stanbon(75887)

Solve (16^x)(2^(x-2))=(1/64)
-----
(2^(4x))(2^(x-2) = 2^-4
----
2^(4x+x-2) = 2^(-4)
---
5x-2 = -4
5x = -2
x = -2/5
=================
Cheers,
Stan H.